OpenStudy (anonymous):

Show all work to simplify 2 over x + 2 over x minus 1 - 2 over x minus 2. Use complete sentences to explain how to simplify this expression

OpenStudy (campbell_st):

so is the question \[\frac{\frac{2}{x + 2}}{x -1} - \frac{2}{x - 2}\]

OpenStudy (anonymous):



OpenStudy (campbell_st):

ok... so you need a common denominator 1st term is \[\frac{2}{x} \times \frac{(x -1)(x -2)}{(x - 1)(x - 2)}\] 2nd term is \[\frac{2}{(x -1)} \times \frac{x(x -2)}{x(x -2)}\] and the last term is \[\frac{2}{x - 2} \times \frac{x(x -1)}{x(x -1)}\] so you will have \[\frac{2(x -1)(x-2) + 2x(x -2) - 2x(x - 1)}{x(x -1)(x-2)}\] so you need to distribute \[\frac{2x^2 - 6x + 4 + 2x^2 - 4x - 2x^2 + 2x}{x^3 - 3x^2 + 2x}\] lastly just collect like terms.

OpenStudy (anonymous):

thank you so much!